if A and B differs in their age by 60 years. After 10 years, the sum of their ages is 100, wats the present age of both
Answer by ankor@dixie-net.com(22740)

Write an equation for each statement, then use elimination
:
if A and B differs in their age by 60 years.
a - b = 60
:
After 10 years, the sum of their ages is 100,
(a+10) + (b+10) = 100
a + b = 100 - 20
a + b = 80
:
a - b = 60
a + b = 80
------------addition eliminates b, find a
2a = 140
a = 70 yrs is A's age
I'll let you find b




wats the present age of both